# Injecting External Knowledge into the Reasoning Process Enhances Retrieval-Augmented Generation

### v1

@@ Abstract.
Retrieval-augmented generation (RAG) has been widely adopted to augment large language models (LLMs) with external knowledge for knowledge-intensive tasks. However, its effectiveness is often undermined by the presence of noisy (i.e., low-quality) retrieved passages. Enhancing LLMs’ robustness to such noise is critical for improving the reliability of RAG systems. Recent advances have equipped LLMs with strong reasoning and self-reflection capabilities, allowing them to identify and correct errors in their reasoning process. Inspired by this ability, we propose Passage Injection—a simple yet effective method that explicitly incorporates retrieved passages into LLMs’ reasoning process, aiming to enhance the model’s ability to recognize and resist noisy passages. We validate Passage Injection under general RAG settings using BM25 as the retriever. Experiments on four reasoning-enhanced LLMs across four factual QA datasets demonstrate that Passage Injection significantly improves overall RAG performance. Further analysis on two noisy retrieval settings-random noise, where the model is provided irrelevant passages, and counterfactual noise, where it is given misleading passages-shows that Passage Injection consistently improves robustness. Controlled experiments confirm that Passage Injection can also effectively leverage helpful passages. These findings suggest that incorporating passages in LLMs’ reasoning process is a promising direction for building more robust RAG systems.
@@
submissionid: 102copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Question answering
@@ 2.1. Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) enhances large language models (LLMs) with a retrieval module that retrieves relevant passages from an external corpus given a query (Lewis et al., 2020; Zamani et al., 2022; Glass et al., 2022; Shi et al., 2024; Zhang et al., 2025b). These retrieved documents are appended to the model’s input context, enabling the LLM to access information beyond its parametric knowledge and improving performance across a wide range of tasks (Lewis et al., 2020; Borgeaud et al., 2022; Izacard et al., 2023). A range of studies has explored how to better integrate these retrieved passages into the generation process, including prompt-level augmentation (Trivedi et al., 2022; Wang et al., 2024), embedding-level fusion (Izacard and Grave, 2020; Dong et al., 2025), and parametric-level adaptation (Su et al., 2025). Our work proposes a new direction by injecting retrieved passages directly into the model’s reasoning process, enabling a tighter coupling between external knowledge and step-by-step reasoning.
@@ 4.1. General RAG Settings
[…]
@@ 5.1. Performance under General RAG
[…]

### v3

@@ Abstract.
Retrieval-augmented generation (RAG) has been widely adopted to augment large language models (LLMs) with external knowledge for knowledge-intensive tasks. However, its effectiveness is often undermined by the presence of noisy (i.e., low-quality) retrieved passages. Enhancing LLMs’ robustness to such noise is critical for improving the reliability of RAG systems. Recent advances have equipped LLMs with strong reasoning and self-reflection capabilities, allowing them to identify and correct errors in their reasoning process. Inspired by this ability, we propose Passage Injection—a simple yet effective method that explicitly incorporates retrieved passages into LLMs’ reasoning process, aiming to enhance the model’s ability to recognize and resist noisy passages. We validate Passage Injection under general RAG settings using BM25 as the retriever. Experiments on four reasoning-enhanced LLMs across four factual QA datasets demonstrate that Passage Injection significantly improves overall RAG performance. Further analysis on two noisy retrieval settings-random noise, where the model is provided irrelevant passages, and counterfactual noise, where it is given misleading passages-shows that Passage Injection consistently improves robustness. Controlled experiments confirm that Passage Injection can also effectively leverage helpful passages. These findings suggest that incorporating passages in LLMs’ reasoning process is a promising direction for building more robust RAG systems.  [^1] The code and data can be found at: https://github.com/Trustworthy-Information-Access/Passage-Injection
@@
submissionid: 102journalyear: 2025copyright: ccconference: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region; December 7–10, 2025; Xi’an, Chinabooktitle: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region (SIGIR-AP 2025), December 7–10, 2025, Xi’an, Chinadoi: 10.1145/3767695.3769515isbn: 979-8-4007-2218-9/2025/12ccs: Information systems Question answeringccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generation
@@ 2.1. Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) enhances large language models (LLMs) with a retrieval module that retrieves relevant passages from an external corpus given a query (Lewis et al., 2020; Zamani et al., 2022; Glass et al., 2022; Shi et al., 2024; Zhang et al., 2025b). These retrieved documents are appended to the model’s input context, enabling the LLM to access information beyond its parametric knowledge and improving performance across a wide range of tasks (Lewis et al., 2020; Borgeaud et al., 2022; Izacard et al., 2023). A range of studies has explored how to better integrate these retrieved passages into the generation process, including prompt-level augmentation (Trivedi et al., 2022; Wang et al., 2024), embedding-level fusion (Izacard and Grave, 2020; Dong et al., 2025), and parametric-level adaptation (Su et al., 2025; Tan et al., 2025). Our work proposes a new direction by injecting retrieved passages directly into the model’s reasoning process, enabling a tighter coupling between external knowledge and step-by-step reasoning.
@@ 4.1. General RAG Settings
To verify whether Passage Injection can improve the performance of RAG systems, we conduct experiments under general RAG settings. For each question, we retrieve the top-$k$ passages from a Wikipedia dump (Ni et al., 2025) using BM25 (Robertson et al., 2009), setting $k$ to 1, 3, and 5 to assess the effect of Passage Injection across different numbers of retrieved passages.
@@ 5.1. Performance under General RAG
Table 1 presents the performance of different RAG methods across four QA benchmarks using top-5 retrieved passages. The main findings are as follows: 1) Passage Injection consistently yields the best performance. Across all models, Passage Injection achieves the highest average F1 scores, demonstrating its effectiveness in general RAG scenarios. Moreover, as shown in Figure 1(b), its performance gains remain stable across different top-$k$ values. 2) Passage Injection brings more gains on multi-hop QA. Compared to its improvements on single-hop PopQA, Passage Injection yields more significant gains on multi-hop datasets. This suggests that Passage Injection is particularly effective for questions requiring complex reasoning. 3) Incorporating passages is crucial. While Instruction Injection yields moderate improvements over Vanilla RAG on most datasets, it remains less effective than Passage Injection. This suggests that explicitly incorporating retrieved passages into the reasoning process is more beneficial than instructions alone. 4) Distilled models benefit more from RAG. Interestingly, DeepSeek-R1-Distill-Qwen-32B underperforms Qwen3-32B in the Direct QA setting but surpasses it under Vanilla RAG. This may be due to the nature of distillation: the model is trained to follow teacher demonstrations that involve long-context reasoning, making it more aligns with the RAG format and better at utilizing retrieved content. 5) Diminished gains on distilled models. While Passage Injection improves all models, the gains are notably smaller for DeepSeek-R1-Distill-Qwen-32B. This may be because its reasoning ability is acquired through supervised fine-tuning from teacher models, rather than developed through its own emergent capabilities. Consequently, the model may be less sensitive to modifications within the <think></think> segment.
@@ 6. Conclusion
Acknowledgements.
This work was funded by the National Natural Science Foundation of China (NSFC) under Grant No. 62302486, the Innovation Funding of ICT CAS under Grant No. E361140, the CAS Special Research Assistant Funding Project, the project under Grants No. JCKY2022130C039, the Strategic Priority Research Program of the CAS under Grants No. XDB0680102, and the NSFC Grant No. 62441229.